A bacteria culture starts with 6500 bacteria. After 2.5 hours,there are 208000 bacteria present. What is the length of the doubling period?..SHOW WORK.
Answer by Boreal(15235)

208000=6500(1+r)^n
divide by 6500
32=(1+r)^n
actually, 2^5=32
so r has to be 1 and n=5
but n=5 means 2.5 hours, so the doubling period is every 30 minutes or half hour.
13K
26K
52K
104K
208K